# Relation between equilibrium quantum phase transitions and dynamical quantum phase transitions in two-band systems

### v1

@@ III The model and results
Here we adopt periodic boundary condition $c_{L+1}=c_{1}$ and take even $L$. After the Fourier transformation $c_{l}^{\dagger}=\frac{1}{\sqrt{L}}\sum_{k}e^{ikl}c_{k}^{\dagger}$, the Hamiltonian in the momentum space can be written as
@@
where $\Psi_{k}=(c_{k},c_{-k}^{\dagger})^{T}$, $\mathcal{K}=\left\{\frac{2\pi m}{L}\mid m=1,2,\dots,\frac{L-1}{2}\right\}$, and $H_{k}$ takes the form of Eq. (5) with the vector components expressed as
$d_{x,k}$ $=-2h-2\cos k,$ (15)
@@
[…]
Choosing $(h,\eta)$ as quench parameters, and quenching them from $(h_{i},\eta_{i})$ to $(h_{f},\eta_{f})$ at time $t=0$, then we get
¯$\bar{\mathcal{L}}_{k}=$ $\frac{[(h_{i}+\cos k)(h_{f}+\cos k)+\eta_{i}\eta_{f}\sin^{2}k]^{2}}{[(h_{i}+\cos k)^{2}+\eta_{i}^{2}\sin^{2}k][(h_{f}+\cos k)^{2}+\eta_{f}^{2}\sin^{2}k]},$ (17)
$\mathcal{F}_{k}^{\text{q}}=$ $\sqrt{\frac{(h_{i}+\cos k)(h_{f}+\cos k)+\eta_{i}\eta_{f}\sin^{2}k}{2\sqrt{[(h_{i}+\cos k)^{2}+\eta_{i}^{2}\sin^{2}k][(h_{f}+\cos k)^{2}+\eta_{f}^{2}\sin^{2}k]}}+\frac{1}{2}}.$ (18)
@@
To exhibit the distribution of post-quench parameters which possess ¯$\bar{\mathcal{L}}_{k}=0$ for $h_{i}=-2$ and $\eta_{i}=0.8$, we plot Fig. 2(a) where cyan solid lines are post-quench parameters which possess ¯$\bar{\mathcal{L}}_{k}=0$ with $k\in\left\{2\pi m/L\mid m=1,2,\dots,(L-1)/2\right\}$ from lighter to darker, here $L=30$. Each post-quench parameter point has as many $k_{c}$-modes which satisfy ¯$\bar{\mathcal{L}}_{k_{c}}=0$ as the number of cyan solid lines that pass through it. According to the number $n_{k_{c}}$ of $k_{c}$-modes owned by the post-qench parameter points, we plot the dynamical phase diagram of DQPTs of the XY model in the thermodynamic limit for $h_{i}=-2$ and $\eta_{i}=0.8$ as shown in Fig. 2(b). Compared with Fig. 1(a), the two additional dashed lines in Fig. 2(b) demonstrate differences between dynamical phase boundaries and equilibrium phase boundaries. Obviously, DQPTs will occur when post-quench parameters are in yellow and orange regions in Fig. 2(b). Since $k_{c}$-modes may exist for post-quench parameters and pre-quench parameters being in the same phase, even quenching within the same equilibrium phase may also lead to the occurrence of DQPTs.
@@
Figure 2: (a) The distribution of $k_{c}$-modes ¯$\bar{\mathcal{L}}_{k_{c}}$ in the post-quench parameter space spanned by $h_{f}$ and $\eta_{f}$. The cyan solid lines are post-quench parameters which possess ¯$\bar{\mathcal{L}}_{k}=0$ with $k\in\left\{2\pi m/L\mid m=1,2,\dots,(L-1)/2\right\}$ from lighter to darker. The purple dotted line denotes $\eta_{f}=-2$. The black dot denotes point $(-2,0.8)$. Here $L=30$. (b) The dynamical phase diagram of DQPTs of the XY model with different numbers $n_{k_{c}}$ of the LE. DQPTs will occur when post-quench parameters are in yellow and orange regions. Here $h_{i}=-2$ and $\eta_{i}=0.8$.
@@
Setting $\eta_{f}=-2$, we plot the density graphic of ¯$\bar{\mathcal{L}}_{k}$ versus $k$ and $h_{f}$ in Fig. 3(a). The cyan solid line represents ¯$\bar{\mathcal{L}}_{k}=0$. As seen from Fig. 3(a), whenever post-quench parameters across dynamical phase boundaries of DQPTs as depicted in Fig. 2(b), there is a change in the number of $k_{c}$-modes but no abrupt change of ¯$\bar{\mathcal{L}}_{k}$. Fig. 3(b) and 3(c) exhibit images of the rate function $\bar{\lambda}=-\frac{1}{L}\sum_{k}\ln\bar{\mathcal{L}}_{k}$ and its derivative $\bar{\lambda}^{\prime}$ with respect to $h_{f}$. The rate function $\bar{\lambda}$ exhibits non-analyticity at all three dynamical phase boundaries. The derivative $\lambda^{\prime}_{T}$ becomes divergent when approaching a dynamical phase boundary from a region with less number of $k_{c}$-modes to a region with more number of $k_{c}$-modes, while it keeps finite at the other side. We cannot distinguish the additional phase boundary of DQPTs from the two boundaries that coincides with equilibrium phase boundaries through the image of the rate function, because the abrupt changes of $k$-modes $\mathcal{F}_{k}^{\text{q}}$ with $k=0$ and $k=\pi$ from 1 to 0 or from 0 to 1 at the two equilibrium phase boundaries are unchanged for ¯$\bar{\mathcal{L}}_{k}$ according to Eq. (11). There is only a change of the number of $k_{c}$-modes but no abrupt change of ¯$\bar{\mathcal{L}}_{k}$ at all three boundaries.
@@
Figure 3: (a) The density graphic of ¯$\bar{\mathcal{L}}_{k}$ versus $k$ and $h_{f}$. The cyan solid line represents ¯$\bar{\mathcal{L}}_{k}=0$. (b) The image of the rate function $\bar{\lambda}$ versus $h_{f}$. (c) The image of the derivative of the rate function $\bar{\lambda}^{\prime}$ versus $h_{f}$. Here $h_{i}=-2$, $\eta_{i}=0.8$, and $\eta_{f}=-2$.
@@
For consistency, we call $k$-modes satisfying $\mathcal{F}_{k}^{\text{q}}=0$ as $k_{0}$-modes and call $k$-modes satisfying $\mathcal{F}_{k}^{\text{q}}=1$ as $k_{1}$-modes in the following. To exhibit how $k$-modes $\mathcal{F}_{k}^{\text{q}}$ change at equilibrium phase boundaries, we display the distribution of $k_{0}$-modes $\mathcal{F}_{k_{0}}^{\text{q}}$ and $k_{1}$-modes $\mathcal{F}_{k_{1}}^{\text{q}}$ in the space of post-quench parameter with the fixed pre-quench parameters $h_{i}=-2$ and $\eta_{i}=0.8$ in Fig. 4(a). The red dashed lines and green dotted lines depict post-quench parameters which possess $\mathcal{F}_{k}^{\text{q}}=1$ and $\mathcal{F}_{k}^{\text{q}}=0$ with $k\in\left\{2\pi m/L\mid m=1,2,\dots,(L-1)/2\right\}$ from lighter to darker, respectively. Here $L=30$. In addition, for $k=0$, we have $\mathcal{F}_{0}^{\text{q}}=1$ for $h_{f}<-1$, and $\mathcal{F}_{0}^{\text{q}}=0$ for $h_{f}>-1$. For $k=\pi$, we have $\mathcal{F}_{\pi}^{\text{q}}=1$ for $h_{f}<1$, and $\mathcal{F}_{\pi}^{\text{q}}=0$ for $h_{f}>1$. It is evident that there are only $k_{1}$-modes but no $k_{0}$-modes for $h_{f}<-1$, there are both $k_{0}$-modes and $k_{1}$-modes for $-1<h_{f}<1$, and there are only $k_{0}$-modes but no $k_{1}$-modes for $h_{f}>1$. From Fig. 4(a), we can conclude that each point on equilibrium phase boundaries corresponds to a $k$-mode $\mathcal{F}_{k}^{\text{q}}$ whose value will suddenly change at that point and smoothly change everywhere else. Setting $h_{i}=-2$ and $\eta_{i}=0.8$, we divide the post-quench parameter space in the thermodynamic limit into four kinds of regions as shown in Fig. 4(b) and 4(c), according to the number $n_{k_{0}}$ of $k_{0}$-modes which satisfy $\mathcal{F}_{k_{0}}=0$ and the number $n_{k_{1}}$ of $k_{1}$-modes which satisfy $\mathcal{F}_{k_{1}}=1$, respectively. Although Fig. 4(b) and 4(c) are a little different from Fig. 1, the overlapping boundaries of Fig. 4(b) and 4(c) match exactly the equilibrium phase boundaries of Fig. 1. It implies both $n_{k_{0}}$ and $n_{k_{1}}$ change at equilibrium phase boundaries. Comparing Fig. 4(b) with Fig. 1, we find that Fig. 4(b) has two extra dashed lines across which there will appear more or less $k_{0}$-modes, whereas abrupt change in each $\mathcal{F}_{k}^{\text{q}}$ only occurs at equilibrium phase boundaries and not here, which we will demonstrate later. Besides, positions of these two extra dashed lines depend on the value of $h_{i}$ and $\eta_{i}$ and thus are not robust.
@@
Figure 4: (a) The distribution of $k_{0}$-modes $\mathcal{F}_{k_{0}}^{\text{q}}$ and $k_{1}$-modes $\mathcal{F}_{k_{1}}^{\text{q}}$ in the post-quench parameter space spanned by $h_{f}$ and $\eta_{f}$. The red dashed lines and green dotted lines depict post-quench parameters which possess $\mathcal{F}_{k}^{\text{q}}=1$ and $\mathcal{F}_{k}^{\text{q}}=0$ with $k\in\left\{2\pi m/L\mid m=1,2,\dots,(L-1)/2\right\}$ from lighter to darker, respectively. The post-quench parameters in dark blue region $h_{f}<-1$ possess both $\mathcal{F}_{0}^{\text{q}}=1$ and $\mathcal{F}_{\pi}^{\text{q}}=1$. The post-quench parameters in light blue region $-1<h_{f}<1$ possess $\mathcal{F}_{0}^{\text{q}}=0$ and $\mathcal{F}_{\pi}^{\text{q}}=1$. The post-quench parameters in faint blue region $h_{f}>1$ possess both $\mathcal{F}_{0}^{\text{q}}=0$ and $\mathcal{F}_{\pi}^{\text{q}}=0$. The purple dotted line denotes $\eta_{f}=-2$. The black dot denotes point $(-2,0.8)$. Here $L=30$. (b) and (c) are regions with different values of $n_{k_{0}}$ and $n_{k_{1}}$, respectively. Solid black lines are original equilibrium phase boundaries, and dashed black lines are additional boundaries. Here $h_{i}=-2$ and $\eta_{i}=0.8$.
@@
Setting $\eta_{f}=-2$, we plot the density graphic of $\mathcal{F}_{k}^{\text{q}}$ versus $k$ and $h_{f}$ in Fig. 5(a). The red dashed lines represent $\mathcal{F}_{k}^{\text{q}}=1$, and the green dotted lines represent $\mathcal{F}_{k}^{\text{q}}=0$. As seen from Fig. 5(a), when the post-quench parameter across the equilibrium phase boundaries $h_{f}=-1$ and $h_{f}=1$, there is a $k$-mode $\mathcal{F}_{k}^{\text{q}}$ abruptly drops from $1$ to 00. When the parameter across the extra boundary in Fig. 4(b), there is only a change in the number of $k_{0}$-modes without any abrupt change in each $\mathcal{F}_{k}^{\text{q}}$. Fig. 5(b) and 5(c) exhibit images of the decay rate function of the quench fidelity $\alpha^{\text{q}}=-\frac{1}{L}\ln\mathcal{F}^{\text{q}}$ and its derivative $\alpha^{\text{q}}$ with respect to $h_{f}$. The decay rate function exhibits non-analyticity at all three boundaries. Nonetheless, the derivative $\alpha^{\text{q}}$ is divergent around the two equilibrium phase boundaries but is finite at one side of the additional boundary. It is because there is only a change of the number of $k_{0}$-modes but no abrupt change of $\mathcal{F}_{k}^{\text{q}}$ at the extra boundary, while both of them change at equilibrium phase boundaries. As a result, the behavior of the derivative $\alpha^{\text{q}}$ can distinguish equilibrium phase boundaries from extra boundaries.
@@
Figure 5: (a) The density graphic of $\mathcal{F}_{k}^{\text{q}}$ versus $k$ and $h_{f}$. The red dashed lines and green dotted lines represent $\mathcal{F}_{k}^{\text{q}}=1$ and $\mathcal{F}_{k}^{\text{q}}=0$, respectively. (b) The image of the decay rate function $\alpha^{\text{q}}$ versus $h_{f}$]. (c) The image of the derivative of the decay rate function $\alpha^{\text{q}}$ versus $h_{f}$. Here $h_{i}=-2$, $\eta_{i}=0.8$ and $\eta_{f}=-2$.
@@
Next we explore the relation between $\mathcal{F}_{k}^{\text{q}}$ and ¯$\bar{\mathcal{L}}_{k}$ of the XY model. According to Eq. (11), we plot the image of ¯$\bar{\mathcal{L}}_{k}$ versus $\mathcal{F}_{k}^{\text{q}}$ in Fig. 6(a). It demonstrates that ¯$\bar{\mathcal{L}}_{k}$ must undergo a zero value as long as $\mathcal{F}_{k}^{\text{q}}$ goes from 00 to $1$. Because $\mathcal{F}_{k}^{\text{q}}$ is a continuous function of $k$, if the fidelity $\mathcal{F}^{\text{q}}$ has a $k_{0}$-mode which satisfies $\mathcal{F}_{k_{0}}^{\text{q}}=0$ and a $k_{1}$-mode which satisfies $\mathcal{F}_{k_{1}}^{\text{q}}=1$, then $\mathcal{F}^{\text{q}}$ must have at least a $k_{c}$mode between $k_{0}$ and $k_{1}$ which satisfies $\mathcal{F}_{k_{c}}^{\text{q}}=\sqrt{2}/2$ so that ¯$\bar{\mathcal{L}}_{k_{c}}=0$. As a consequence, a sufficient condition for the occurrence of DQPTs is the quench fidelity $\mathcal{F}^{\text{q}}$ has at least a $k_{0}$-mode and a $k_{1}$-mode simultaneously. To make it more intuitional, we take the four post-quench parameter points denoted by purple circle on the purple dashed line in Fig. 4(a) as cases, and plot images of $\mathcal{F}_{k}^{\text{q}}$ and ¯$\bar{\mathcal{L}}_{k}$ versus $k$ in Fig. 6 (b)-(e). The dark blue and light blue solid lines represent $\mathcal{F}_{k}^{\text{q}}$ and ¯$\bar{\mathcal{L}}_{k}$, respectively. The blue horizontal dashed lines represent 2/$\sqrt{2}/2$, and the black vertical dashed lines denote corresponding $k_{c}$. Setting $h_{i}=-2$ and $\eta_{i}=0.8$, Fig. 6(b) and 6(e) show that the quench fidelity only have $k_{1}$-modes or $k_{0}$-modes and the images of $\mathcal{F}_{k}^{\text{q}}$ do not cross the dashed line, thus there is no $k_{c}$-mode satisfies ¯$\bar{\mathcal{L}}_{k_{c}}=0$, which means no DQPT will happen. Fig. 6(d) shows the quench fidelity has both a $k_{0}$-mode and a $k_{1}$-mode, and the image of $\mathcal{F}_{k}^{\text{q}}$ certainly cross the horizontal dashed line, thus ¯$\bar{\mathcal{L}}_{k}$ becomes 0 at the cross point $k_{c}$ and DQPTs can occur. Besides, even if the quench fidelity only have $k_{1}$-modes and have no $k_{0}$-mode, the image of $\mathcal{F}_{k}^{\text{q}}$ can still cross the dashed line, which leads to the occurrence of DQPTs as shown in Fig. 6(c). This is consistent with Fig. 2(b) and Fig. 4(a): The quench fidelity has both $k_{0}$-modes and $k_{1}$-modes for $-1<h_{f}<1$, thus DQPTs must occur when post-quench parameters are in this region. While for $h_{f}<-1$ and $h_{f}>1$, the quench fidelity has either only $k_{1}$-modes or only $k_{0}$-modes, thus DQPTs may or may not occur in these regions. In a nutshell, since the necessary and sufficient condition for DQPTs is the quench fidelity of the pre-quench system and the post-quench system has at least a $k_{c}$-mode which satisfies $\mathcal{F}_{k_{c}}^{\text{q}}=\sqrt{2}/2$, whereas the necessary and sufficient condition for crossing the EQPT point is the quench fidelity has at least a $k_{0}$-mode which satisfies $\mathcal{F}_{k_{0}}^{\text{q}}=0$, crossing the EQPT point is neither sufficient nor necessary for the occurrence of DQPTs.

### v2

@@ III The model and results
Here we adopt the periodic boundary condition $c_{L+1}=c_{1}$ and take even $L$. After the Fourier transformation $c_{l}^{\dagger}=\frac{1}{\sqrt{L}}\sum_{k}e^{ikl}c_{k}^{\dagger}$, the Hamiltonian in the momentum space can be written as
@@
where $\Psi_{k}=(c_{k},c_{-k}^{\dagger})^{T}$, $\mathcal{K}=\left\{\frac{2\pi m}{L}\mid m=1,2,\dots,\frac{L-2}{2}\right\}$, and $H_{k}$ takes the form of Eq. (5) with the vector components expressed as
$d_{z,k}$ $=-2h-2\cos k,$ (17)
@@
$d_{x,k}$ $=d_{0,k}=0.$ (19)
The ground state phase diagram of the XY model is shown in Fig. 1. There are three equilibrium phase boundaries: the critical line h=1h=-1 where the energy gap closes at $k=0$, the critical line $h=1$ where the energy gap closes at $k=\pi$, and the critical line $\eta=0$ for 1<h<1-1<h<1 where the energy gap closes at $k=\pm\arccos(-h)$. The winding number $\nu$ is defined by $\nu=\frac{1}{2\pi}\int_{-\pi}^{\pi}dk\frac{d_{x,k}\partial_{k}d_{y,k}-d_{y,k}\partial_{k}d_{x,k}}{d_{x,k}^{2}+d_{y,k}^{2}}$.
Choose $\gamma=(h,\eta)$ as quench parameters, and quench them from $\gamma_{i}=(h_{i},\eta_{i})$ to $\gamma_{f}=(h_{f},\eta_{f})$ at time $t=0$. Substituting the expression of $\mathbf{d}_{k}(\gamma_{i/f})$ given by Eqs. (17)-(19) into Eq. (10) and Eq. (11), we can get the expressions of $\mathcal{F}_{k}^{\text{q}}$ and $\bar{\mathcal{L}}_{k}$, which are explicitly given by
$\mathcal{F}_{k}^{\text{q}}=$ $\sqrt{\frac{(h_{i}+\cos k)(h_{f}+\cos k)+\eta_{i}\eta_{f}\sin^{2}k}{2\sqrt{[(h_{i}+\cos k)^{2}+\eta_{i}^{2}\sin^{2}k][(h_{f}+\cos k)^{2}+\eta_{f}^{2}\sin^{2}k]}}+\frac{1}{2}},$ (20)
$\bar{\mathcal{L}}_{k}=$ $\frac{[(h_{i}+\cos k)(h_{f}+\cos k)+\eta_{i}\eta_{f}\sin^{2}k]^{2}}{[(h_{i}+\cos k)^{2}+\eta_{i}^{2}\sin^{2}k][(h_{f}+\cos k)^{2}+\eta_{f}^{2}\sin^{2}k]}.$ (21)
@@
To clearly exhibit the distribution of postquench parameters which possess $\bar{\mathcal{L}}_{k}=0$ for $h_{i}=-2$ and $\eta_{i}=0.8$, we plot Fig. 2(a) according to Eq. (21) for the case $L=30$, where the cyan solid lines are postquench parameters which possess $\bar{\mathcal{L}}_{k}=0$ with $k=2\pi/L,4\pi/L,\dots,(L-2)\pi/L$ from lighter to darker, respectively. Each postquench parameter point has as many $k_{c}$-modes which satisfy $\bar{\mathcal{L}}_{k_{c}}=0$ as the number of cyan solid lines that pass through it. As the system size increases, the cyan solid lines become denser and denser. According to the number $n_{k_{c}}$ of $k_{c}$-modes owned by the postquench parameter points, we obtain the dynamical phase diagram of DQPTs of the XY model in the thermodynamic limit for $h_{i}=-2$ and $\eta_{i}=0.8$ as shown in Fig. 2(b). Compared with Fig. 1, the two additional dashed lines in Fig. 2(b) demonstrate differences between dynamical phase boundaries and equilibrium phase boundaries. Obviously, DQPTs will occur when postquench parameters are in yellow and orange regions in Fig. 2(b). Since $k_{c}$-modes may exist for postquench parameters and prequench parameters being in the same phase, even quenching within the same equilibrium phase may also lead to the occurrence of DQPTs.
@@
Figure 2: (a) The distribution of $k_{c}$-modes $\bar{\mathcal{L}}_{k_{c}}$ in the postquench parameter space spanned by $h_{f}$ and $\eta_{f}$. The cyan solid lines are postquench parameters which possess $\bar{\mathcal{L}}_{k}=0$ with $k=2\pi/L,4\pi/L,\dots,(L-2)\pi/L$ from lighter to darker, respectively. The purple dotted line denotes $\eta_{f}=-2$. The black dot denotes point (2,0.8)(-2,0.8). Here $L=30$. (b) The dynamical phase diagram of DQPTs of the XY model with different numbers $n_{k_{c}}$ of the LE. DQPTs will occur when postquench parameters are in yellow and orange regions. Here $h_{i}=-2$ and $\eta_{i}=0.8$.
@@
Setting $\eta_{f}=-2$, we plot the density graphic of $\bar{\mathcal{L}}_{k}$ versus $k$ and $h_{f}$ in Fig. 3(a). The cyan solid line represents $\bar{\mathcal{L}}_{k}=0$. As seen from Fig. 3(a), whenever postquench parameters cross dynamical phase boundaries of DQPTs as depicted in Fig. 2(b), there is a change in the number of $k_{c}$-modes but no abrupt change of $\bar{\mathcal{L}}_{k}$. Figures 3(b) and 3(c) exhibit images of the rate function $\bar{\lambda}=-\frac{1}{L}\sum_{k}\ln\bar{\mathcal{L}}_{k}$ and its derivative $\bar{\lambda}^{\prime}$ with respect to $h_{f}$. The rate function $\bar{\lambda}$ exhibits nonanalyticity at all three dynamical phase boundaries. The derivative $\bar{\lambda}^{\prime}$ becomes divergent when approaching a dynamical phase boundary from a region with fewer $k_{c}$-modes to a region with more $k_{c}$-modes, while it remains finite at the other side. We cannot distinguish the additional phase boundary of DQPTs from the two boundaries that coincide with equilibrium phase boundaries through the image of the rate function, because the abrupt changes of $k$-modes $\mathcal{F}_{k}^{\text{q}}$ with $k=0$ and $k=\pi$ from 1 to 0 or from 0 to 1 at the two equilibrium phase boundaries are unchanged for $\bar{\mathcal{L}}_{k}$ according to Eq. (11). There is only a change of the number of $k_{c}$-modes but no abrupt change of $\bar{\mathcal{L}}_{k}$ at all three boundaries.
@@
Figure 3: (a) The density graphic of $\bar{\mathcal{L}}_{k}$ versus $k$ and $h_{f}$. The cyan solid line represents $\bar{\mathcal{L}}_{k}=0$. (b) The image of the rate function $\bar{\lambda}$ versus $h_{f}$. (c) The image of the derivative of the rate function $\bar{\lambda}^{\prime}$ versus $h_{f}$. Here $h_{i}=-2$, $\eta_{i}=0.8$, and $\eta_{f}=-2$.
@@
For consistency, we call $k$-modes satisfying $\mathcal{F}_{k}^{\text{q}}=0$ as $k_{0}$-modes and call $k$-modes satisfying $\mathcal{F}_{k}^{\text{q}}=1$ as $k_{1}$-modes in the following. To clearly exhibit how $k$-modes $\mathcal{F}_{k}^{\text{q}}$ abruptly change at equilibrium phase boundaries, we display the distribution of $k_{0}$-modes $\mathcal{F}_{k_{0}}^{\text{q}}$ and $k_{1}$-modes $\mathcal{F}_{k_{1}}^{\text{q}}$ according to Eq. (20) in the space of postquench parameters in Fig. 4(a) with system size $L=30$, where fixed prequench parameters $h_{i}=-2$ and $\eta_{i}=0.8$. The red dashed lines and green dotted lines depict postquench parameters which possess $\mathcal{F}_{k}^{\text{q}}=1$ and $\mathcal{F}_{k}^{\text{q}}=0$ with $k=2\pi/L,4\pi/L,\dots,(L-2)\pi/L$ from lighter to darker, respectively. In addition, for $k=0$, $\mathcal{F}_{0}^{\text{q}}=1$ for $h_{f}<-1$, and $\mathcal{F}_{0}^{\text{q}}=0$ for $h_{f}>-1$. For $k=\pi$, $\mathcal{F}_{\pi}^{\text{q}}=1$ for $h_{f}<1$, and $\mathcal{F}_{\pi}^{\text{q}}=0$ for $h_{f}>1$. It is evident that there are only $k_{1}$-modes but no $k_{0}$-modes for $h_{f}<-1$, there are both $k_{0}$-modes and $k_{1}$-modes for $1<h_{f}<1$, and there are only $k_{0}$-modes but no $k_{1}$-modes for $h_{f}>1$. From Fig. 4(a), we can conclude that each point on equilibrium phase boundaries corresponds to a $k$-mode $\mathcal{F}_{k}^{\text{q}}$ whose value will suddenly change at that point and smoothly change everywhere else. As the system size increases, the red dashed lines and green dotted lines become denser and denser. Setting $h_{i}=-2$ and $\eta_{i}=0.8$, we divide the postquench parameter space in the thermodynamic limit into four kinds of regions as shown in Figs. 4(b) and 4(c), according to the number $n_{k_{0}}$ of $k_{0}$-modes which satisfy $\mathcal{F}^{\text{q}}_{k_{0}}=0$ and the number $n_{k_{1}}$ of $k_{1}$-modes which satisfy $\mathcal{F}^{\text{q}}_{k_{1}}=1$, respectively. Although Figs. 4(b) and 4(c) are a little different from Fig. 1, the overlapping boundaries of Figs. 4(b) and 4(c) match exactly the equilibrium phase boundaries of Fig. 1. This implies that both $n_{k_{0}}$ and $n_{k_{1}}$ change at equilibrium phase boundaries. Comparing Fig. 4(b) with Fig. 1, we find that Fig. 4(b) has two extra dashed lines across which there will appear more or fewer $k_{0}$-modes, whereas abrupt change in each $\mathcal{F}_{k}^{\text{q}}$ only occurs at equilibrium phase boundaries and not here, which we will demonstrate later. Besides, positions of these two extra dashed lines depend on the value of $h_{i}$ and $\eta_{i}$ and thus are not robust.
@@
Figure 4: (a) The distribution of $k_{0}$-modes $\mathcal{F}_{k_{0}}^{\text{q}}$ and $k_{1}$-modes $\mathcal{F}_{k_{1}}^{\text{q}}$ in the postquench parameter space spanned by $h_{f}$ and $\eta_{f}$. The red dashed lines and green dotted lines depict postquench parameters which possess $\mathcal{F}_{k}^{\text{q}}=1$ and $\mathcal{F}_{k}^{\text{q}}=0$ with $k=2\pi/L,4\pi/L,\dots,(L-2)\pi/L$ from lighter to darker, respectively. The postquench parameters in dark blue region $h_{f}<-1$ possess both $\mathcal{F}_{0}^{\text{q}}=1$ and $\mathcal{F}_{\pi}^{\text{q}}=1$. The postquench parameters in light blue region $1<h_{f}<1$ possess $\mathcal{F}_{0}^{\text{q}}=0$ and $\mathcal{F}_{\pi}^{\text{q}}=1$. The postquench parameters in faint blue region $h_{f}>1$ possess both $\mathcal{F}_{0}^{\text{q}}=0$ and $\mathcal{F}_{\pi}^{\text{q}}=0$. The purple dotted line denotes $\eta_{f}=-2$. The black dot denotes point (2,0.8)(-2,0.8). Here $L=30$. (b) and (c) are regions with different values of $n_{k_{0}}$ and $n_{k_{1}}$, respectively. Solid black lines are original equilibrium phase boundaries, and dashed black lines are additional boundaries. Here $h_{i}=-2$ and $\eta_{i}=0.8$.
@@
Setting $\eta_{f}=-2$, we plot the density graphic of $\mathcal{F}_{k}^{\text{q}}$ versus $k$ and $h_{f}$ in Fig. 5(a). The red dashed lines represent $\mathcal{F}_{k}^{\text{q}}=1$, and the green dotted lines represent $\mathcal{F}_{k}^{\text{q}}=0$. As seen from Fig. 5(a), when the postquench parameter crosses the equilibrium phase boundaries $h_{f}=-1$ and $h_{f}=1$, there is a $k$-mode $\mathcal{F}_{k}^{\text{q}}$ abruptly dropping from $1$ to 0. When the parameter crosses the extra boundary in Fig. 4(b), there is only a change in the number of $k_{0}$-modes without any abrupt change in each $\mathcal{F}_{k}^{\text{q}}$. Figures 5(b) and 5(c) exhibit images of the decay rate function of the quench fidelity $\alpha^{\text{q}}=-\frac{1}{L}\ln\mathcal{F}^{\text{q}}$ and its derivative $\alpha^{\text{q}}$ with respect to $h_{f}$. The decay rate function exhibits nonanalyticity at all three boundaries. Nonetheless, the derivative $\alpha^{\text{q}}$ is divergent around the two equilibrium phase boundaries but is finite at one side of the additional boundary. It is because there is only a change of the number of $k_{0}$-modes but no abrupt change of $\mathcal{F}_{k}^{\text{q}}$ at the extra boundary, while both of them change at equilibrium phase boundaries. As a result, the behavior of the derivative $\alpha^{\text{q}}$ can distinguish equilibrium phase boundaries from extra boundaries.
@@
Figure 5: (a) The density graphic of $\mathcal{F}_{k}^{\text{q}}$ versus $k$ and $h_{f}$. The red dashed lines and green dotted lines represent $\mathcal{F}_{k}^{\text{q}}=1$ and $\mathcal{F}_{k}^{\text{q}}=0$, respectively. (b) The image of the decay rate function $\alpha^{\text{q}}$ versus $h_{f}$. (c) The image of the derivative of the decay rate function $\alpha^{\text{q}}$ versus $h_{f}$. Here $h_{i}=-2$, $\eta_{i}=0.8$, and $\eta_{f}=-2$.
@@
Next we explore the relation between $\mathcal{F}_{k}^{\text{q}}$ and $\bar{\mathcal{L}}_{k}$ of the XY model. According to Eq. (11), we plot the image of $\bar{\mathcal{L}}_{k}$ versus $\mathcal{F}_{k}^{\text{q}}$ in Fig. 6(a). It demonstrates that $\bar{\mathcal{L}}_{k}$ must undergo a zero value as long as $\mathcal{F}_{k}^{\text{q}}$ goes from 0 to $1$. Because $\mathcal{F}_{k}^{\text{q}}$ is a continuous function of $k$, if $\mathcal{F}^{\text{q}}$ has a $k_{0}$-mode satisfying $\mathcal{F}_{k_{0}}^{\text{q}}=0$ and a $k_{1}$-mode satisfying $\mathcal{F}_{k_{1}}^{\text{q}}=1$, then $\mathcal{F}^{\text{q}}$ must have at least a $k_{c}$ -mode between $k_{0}$ and $k_{1}$ which satisfies $\mathcal{F}_{k_{c}}^{\text{q}}=\sqrt{2}/2$ so that $\bar{\mathcal{L}}_{k_{c}}=0$. As a consequence, a sufficient condition for the occurrence of DQPTs is that the quench fidelity $\mathcal{F}^{\text{q}}$ has at least a $k_{0}$-mode and a $k_{1}$-mode simultaneously. To make it more intuitional, we take the four postquench parameter points denoted by purple circles on the purple dashed line in Fig. 4(a) as four cases, and plot images of $\mathcal{F}_{k}^{\text{q}}$ and $\bar{\mathcal{L}}_{k}$ versus $k$ in Figs. 6(b)–6(e), respectively. The dark blue and light blue solid lines represent $\mathcal{F}_{k}^{\text{q}}$ and $\bar{\mathcal{L}}_{k}$, respectively. The blue horizontal dashed lines represent $\sqrt{2}/2$, and the black vertical dashed lines denote the corresponding $k_{c}$. Setting $h_{i}=-2$ and $\eta_{i}=0.8$, Figs. 6(b) and 6(e) show that the quench fidelity only has $k_{1}$-modes and $k_{0}$-modes, respectively, and the images of $\mathcal{F}_{k}^{\text{q}}$ do not cross the horizontal dashed line; thus there is no $k_{c}$-mode satisfying $\bar{\mathcal{L}}_{k_{c}}=0$, which means no DQPT will happen. Figure 6(d) shows that the quench fidelity has both a $k_{0}$-mode and a $k_{1}$-mode, and the image of $\mathcal{F}_{k}^{\text{q}}$ certainly crosses the horizontal dashed line; thus $\bar{\mathcal{L}}_{k}$ becomes 0 at the cross point $k_{c}$ and DQPTs can occur. Besides, even if the quench fidelity only has $k_{1}$-modes and has no $k_{0}$-mode, the image of $\mathcal{F}_{k}^{\text{q}}$ can still cross the horizontal dashed line, which leads to the occurrence of DQPTs as shown in Fig. 6(c). This is consistent with Fig. 2 and Fig. 4: The quench fidelity has both $k_{0}$-modes and $k_{1}$-modes for $1<h_{f}<1$, and thus DQPTs must occur when postquench parameters are in this region, while for $h_{f}<-1$ and $h_{f}>1$, the quench fidelity has either only $k_{1}$-modes or only $k_{0}$-modes, and thus DQPTs may or may not occur in these regions. In a nutshell, since a necessary and sufficient condition for DQPTs is that the quench fidelity of the prequench system and the postquench system has at least a $k_{c}$-mode which satisfies $\mathcal{F}_{k_{c}}^{\text{q}}=\sqrt{2}/2$, whereas the necessary and sufficient condition for crossing the EQPT point is the quench fidelity having at least a $k_{0}$-mode which satisfies $\mathcal{F}_{k_{0}}^{\text{q}}=0$, crossing the EQPT point is neither sufficient nor necessary for the occurrence of DQPTs.